# Privacy-Preserving Federated Learning Scheme with Mitigating Model Poisoning Attacks: Vulnerabilities and Countermeasures

### v1

@@ 4.3.1 Poisonous Gradient Baseline Finding
In the procedure of poisonous gradient baseline finding, $S_{1}$ interacts with $S_{2}$ to measure the secure cosine similarity $\cos_{i\Sigma}\leftarrow\mathbf{SecCos}([\![g_{i}^{(t)}]\!],[\![g_{\Sigma}^{(t-1)}]\!])$ for $i\in[1,n]$, so as to find the baseline of poisonous gradient $[\![g_{*}^{(t)}]\!]\in\{[\![g_{i}^{(t)}]\!]\}_{i\in[1,n]}$ with the lowest cosine similarity. Specifically, in this procedure, $S_{1}$ selects $l$ random noises $\{r_{i1},\cdots,r_{il}\}$ and masks $[\![g_{i}^{(t)}]\!]=\{[\![x_{i1}]\!],\cdots,[\![x_{il}]\!]\}$ and $[\![g_{\Sigma}^{(t-1)}]\!]=\{[\![y_{1}]\!],\cdots,[\![y_{l}]\!]\}$ as $[\![\bar{g}_{i}^{(t)}]\!]$ and $[\![\bar{g}_{\Sigma,i}^{(t)}]\!]$ using Eq. (1) and Eq. (2), respectively. Then $S_{1}$ partially decrypts $[\![\bar{g}_{i}^{(t)}]\!]$ and $[\![\bar{g}_{\Sigma,i}^{(t)}]\!]$ as $[\bar{g}_{i}^{(t)}]_{1}$ and $[\bar{g}_{\Sigma,i}^{(t)}]_{1}$ and sends $[\![\bar{g}_{i}^{(t)}]\!],[\![\bar{g}_{\Sigma,i}^{(t)}]\!],[\bar{g}_{i}^{(t)}]_{1},[\bar{g}_{\Sigma,i}^{(t)}]_{1}$ to $S_{2}$.
$\small\left(\begin{array}[]{c}[\![\bar{g}_{1}^{(t)}]\!]\\ \vdots\\ {[\![\bar{g}_{i}^{(t)}]\!]}\\ \vdots\\ {[\![\bar{g}_{n}^{(t)}]\!]}\\ \end{array}\right)=\left(\begin{array}[]{ccc}[\![x_{11}]\!]\cdot[\![r_{11}]\!]&\cdots&[\![x_{1l}]\!]\cdot[\![r_{1l}]\!]\\ \vdots&\cdot&\vdots\\ {[\![x_{i1}]\!]\cdot[\![r_{i1}]\!]}&\cdots&[\![x_{il}]\!]\cdot[\![r_{il}]\!]\\ \vdots&\cdot&\vdots\\ {[\![x_{n1}]\!]\cdot[\![r_{n1}]\!]}&\cdots&[\![x_{nl}]\!]\cdot[\![r_{nl}]\!]\\ \end{array}\right)$ (1)
@@
$\small\left(\begin{array}[]{c}[\![\bar{g}_{\Sigma,1}^{(t-1)}]\!]\\ \vdots\\ {[\![\bar{g}_{\Sigma,i}^{(t-1)}]\!]}\\ \vdots\\ {[\![\bar{g}_{\Sigma,n}^{(t-1)}]\!]}\\ \end{array}\right)=\left(\begin{array}[]{ccccc}[\![y_{1}]\!]\cdot[\![r_{11}]\!]&\cdots&[\![y_{l}]\!]\cdot[\![r_{1l}]\!]\\ \vdots&\cdot&\vdots\\ {[\![y_{1}]\!]\cdot[\![r_{i1}]\!]}&\cdots&[\![y_{l}]\!]\cdot[\![r_{il}]\!]\\ \vdots&\cdot&\vdots\\ {[\![y_{1}]\!]\cdot[\![r_{n1}]\!]}&\cdots&[\![y_{l}]\!]\cdot[\![r_{nl}]\!]\\ \end{array}\right)$ (2)
Privacy Leakage: Once receiving the ciphertexts $[\![\bar{g}_{i}^{(t)}]\!],[\![\bar{g}_{\Sigma,i}^{(t)}]\!],[\bar{g}_{i}^{(t)}]_{1},[\bar{g}_{\Sigma,i}^{(t)}]_{1}$ from $S_{1}$, the server $S_{2}$ calls $\mathbf{PHE.PartDec}$ and $\mathbf{PHE.FullDec}$ to obtain $\bar{g}_{i}^{(t)}$ and $\bar{g}_{\Sigma,i}^{(t-1)}$ as Eq. (3) and Eq. (4), respectively.
$\small\left(\begin{array}[]{c}\bar{g}_{1}^{(t)}\\ \vdots\\ {\bar{g}_{i}^{(t)}}\\ \vdots\\ {\bar{g}_{n}^{(t)}}\\ \end{array}\right)=\left(\begin{array}[]{ccc}x_{11}+r_{11}&\cdots&x_{1l}+r_{1l}\\ \vdots&\cdot&\vdots\\ {x_{i1}+r_{i1}}&\cdots&x_{il}+r_{il}\\ \vdots&\cdot&\vdots\\ {x_{n1}+r_{n1}}&\cdots&x_{nl}+r_{nl}\\ \end{array}\right)$ (3)
@@
$\small\left(\begin{array}[]{c}\bar{g}_{\Sigma,1}^{(t-1)}\\ \vdots\\ {\bar{g}_{\Sigma,i}^{(t-1)}}\\ \vdots\\ {\bar{g}_{\Sigma,n}^{(t-1)}}\\ \end{array}\right)=\left(\begin{array}[]{ccc}y_{1}+r_{11}&\cdots&y_{l}+r_{1l}\\ \vdots&\cdot&\vdots\\ {y_{1}+r_{i1}}&\cdots&y_{l}+r_{il}\\ \vdots&\cdot&\vdots\\ {y_{1}+r_{n1}}&\cdots&y_{l}+r_{nl}\\ \end{array}\right)$ (4)
Then, $S_{2}$ can learn the difference between two arbitrary local gradients. The difference between the local gradients of the $i$th and $j$th users is calculated as
$\begin{split}d_{ij}&=g_{i}^{(t)}-g_{j}^{(t)}=\{x_{i1}-x_{j1},\cdots,x_{il}-x_{jl}\}\\ &=\left(\bar{g}_{i}^{(t)}-\bar{g}_{j}^{(t)}\right)-\left(\bar{g}_{\Sigma,i}^{(t-1)}-\bar{g}_{\Sigma,j}^{(t-1)}\right).\end{split}$ (5)
@@
Besides, $S_{2}$ can learn the difference between the last-iteration aggregated gradient $g_{\Sigma}^{(t-1)}$ and an arbitrary local gradient $g_{i}^{(t)}$:
@@
4.3.2 Byzantine-Tolerance Aggregation
$S_{1}$ interacts with $S_{2}$ to implement the secure cosine similarity measurement $\mathbf{SecCos}([\![g_{i}^{(t)}]\!],[\![g_{*}^{(t)}]\!])$ for $i=1,\cdots,n$, so as to obtain the similarities of all local gradients related to the poisonous gradient baseline $[\![g_{*}^{(t)}]\!]$. Specifically, in this procedure, $S_{1}$ selects $l$ random noises $\{\gamma_{i1},\cdots,\gamma_{il}\}$ and masks $[\![g_{i}^{(t)}]\!]=\{[\![x_{i1}]\!],\cdots,[\![x_{il}]\!]\}$ and $[\![g_{*}^{(t)}]\!]=\{[\![z_{1}]\!],\cdots,[\![z_{l}]\!]\}$ as $[\![\tilde{g}_{i}^{(t)}]\!]$ and $[\![\tilde{g}_{*,i}^{(t)}]\!]$ using Eq. (8) and Eq. (9), respectively. Then $S_{1}$ partially decrypts $[\![\tilde{g}_{i}^{(t)}]\!]$ and $[\![\tilde{g}_{*,i}^{(t)}]\!]$ as $[\tilde{g}_{i}^{(t)}]_{1}$ and $[\tilde{g}_{*,i}^{(t)}]_{1}$ and sends $[\![\tilde{g}_{i}^{(t)}]\!],[\![\tilde{g}_{*,i}^{(t)}]\!],[\tilde{g}_{i}^{(t)}]_{1},[\tilde{g}_{*,i}^{(t)}]_{1}$ to $S_{2}$.
$\small\left(\begin{array}[]{c}[\![\tilde{g}_{1}^{(t)}]\!]\\ \vdots\\ {[\![\tilde{g}_{i}^{(t)}]\!]}\\ \vdots\\ {[\![\tilde{g}_{n}^{(t)}]\!]}\\ \end{array}\right)=\left(\begin{array}[]{ccc}[\![x_{11}]\!]\cdot[\![\gamma_{11}]\!]&\cdots&[\![x_{1l}]\!]\cdot[\![\gamma_{1l}]\!]\\ \vdots&\cdot&\vdots\\ {[\![x_{i1}]\!]\cdot[\![\gamma_{i1}]\!]}&\cdots&[\![x_{il}]\!]\cdot[\![\gamma_{il}]\!]\\ \vdots&\cdot&\vdots\\ {[\![x_{n1}]\!]\cdot[\![\gamma_{n1}]\!]}&\cdots&[\![x_{nl}]\!]\cdot[\![\gamma_{nl}]\!]\\ \end{array}\right)$ (8)
$\small\left(\begin{array}[]{c}[\![\tilde{g}_{*,1}^{(t-1)}]\!]\\ \vdots\\ {[\![\tilde{g}_{*,i}^{(t-1)}]\!]}\\ \vdots\\ {[\![\tilde{g}_{*,n}^{(t-1)}]\!]}\\ \end{array}\right)=\left(\begin{array}[]{ccccc}[\![z_{1}]\!]\cdot[\![\gamma_{11}]\!]&\cdots&[\![z_{l}]\!]\cdot[\![\gamma_{1l}]\!]\\ \vdots&\cdot&\vdots\\ {[\![z_{1}]\!]\cdot[\![\gamma_{i1}]\!]}&\cdots&[\![z_{l}]\!]\cdot[\![\gamma_{il}]\!]\\ \vdots&\cdot&\vdots\\ {[\![z_{1}]\!]\cdot[\![\gamma_{n1}]\!]}&\cdots&[\![z_{l}]\!]\cdot[\![\gamma_{nl}]\!]\\ \end{array}\right)$ (9)
Privacy Leakage: Upon obtaining the ciphertexts $[\![\tilde{g}_{i}^{(t)}]\!],[\![\tilde{g}_{*,i}^{(t)}]\!],[\tilde{g}_{i}^{(t)}]_{1},[\tilde{g}_{*,i}^{(t)}]_{1}$ from $S_{1}$, the server $S_{2}$ calls $\mathbf{PHE.PartDec}$ and $\mathbf{PHE.FullDec}$ to obtain $\tilde{g}_{i}^{(t)}$ and $\tilde{g}_{*,i}^{(t)}$ as Eq. (10) and Eq. (11), respectively.
$\small\left(\begin{array}[]{c}\tilde{g}_{1}^{(t)}\\ \vdots\\ {\tilde{g}_{i}^{(t)}}\\ \vdots\\ {\tilde{g}_{n}^{(t)}}\\ \end{array}\right)=\left(\begin{array}[]{ccc}x_{11}+\gamma_{11}&\cdots&x_{1l}+\gamma_{1l}\\ \vdots&\cdot&\vdots\\ {x_{i1}+\gamma_{i1}}&\cdots&x_{il}+\gamma_{il}\\ \vdots&\cdot&\vdots\\ {x_{n1}+\gamma_{n1}}&\cdots&x_{nl}+\gamma_{nl}\\ \end{array}\right)$ (10)
$\vspace*{-0.5\baselineskip}\small\left(\begin{array}[]{c}\tilde{g}_{*,1}^{(t)}\\ \vdots\\ {\tilde{g}_{*,i}^{(t)}}\\ \vdots\\ {\tilde{g}_{*,n}^{(t)}}\\ \end{array}\right)=\left(\begin{array}[]{ccc}z_{1}+\gamma_{11}&\cdots&z_{l}+\gamma_{1l}\\ \vdots&\cdot&\vdots\\ {z_{1}+\gamma_{i1}}&\cdots&z_{l}+\gamma_{il}\\ \vdots&\cdot&\vdots\\ {z_{1}+\gamma_{n1}}&\cdots&z_{l}+\gamma_{nl}\\ \end{array}\right)$ (11)
Since $\tilde{g}_{*}^{(t)}$ is the gradient of a certain user, i.e., $\tilde{g}_{*}^{(t)}\in\{\tilde{g}_{1}^{(t)},\cdots,\tilde{g}_{n}^{(t)}\}$, $S_{2}$ can learn which user’s gradient is the poisonous gradient baseline by comparing $\tilde{g}_{*,i}^{(t)}$ and $\tilde{g}_{i}^{(t)}$. That is, if $\tilde{g}_{*,k}^{(t)}=\tilde{g}_{k}^{(t)}$, then $S_{2}$ can determine that the gradient of the $k$th user is the poisonous gradient baseline which has the lowest cosine similarity in the $t$th iteration of the model. This, to some extent, leaks the $k$th user’s personal information. Besides, similar to the analyses in the poisonous gradient baseline finding, in Byzantine-tolerance aggregation, $S_{2}$ can also learn the difference between arbitrary two local gradients from Eq. (10) and Eq. (11). Therefore, $S_{2}$ knows the differences between the gradient of an arbitrary user $i$ and the poisonous gradient baseline, which is
$\begin{split}d_{i*}&=g_{i}^{(t)}-g_{*}^{(t)}=\{x_{i1}-z_{1},\cdots,x_{il}-z_{l}\}\\ &=\left(\tilde{g}_{i}^{(t)}-\tilde{g}_{l}^{(t)}\right)-\left(\tilde{g}_{*,i}^{(t)}-\tilde{g}_{*,l}^{(t)}\right).\end{split}$ (12)
Thus, $S_{2}$ knows a “shifted” distribution of all user gradients related to the poisonous gradient baseline through Eq. (13).
$\small\left(\begin{array}[]{c}g_{1}^{(t)}\\ \vdots\\ {g_{i}^{(t)}}\\ \vdots\\ {g_{n}^{(t)}}\\ \end{array}\right)=\left(\begin{array}[]{c}d_{1*}+g_{*}^{(t)}\\ \vdots\\ {d_{i*}+g_{*}^{(t)}}\\ \vdots\\ {d_{n*}+g_{*}^{(t)}}\\ \end{array}\right).$ (13)
In addition, $S_{2}$ can learn the difference between the poisonous gradient baseline and the last-iteration aggregated gradient:
$\begin{split}d_{*\Sigma}&=g_{*}^{(t)}-g_{\Sigma}^{(t-1)}=\{z_{1}-y_{1},\cdots,z_{l}-y_{l}\}\\ &=\left(\tilde{g}_{*,i}^{(t)}-\bar{g}_{\Sigma,i}^{(t-1)}\right)-\left(\tilde{g}_{i}^{(t)}-\bar{g}_{i}^{(t)}\right).\end{split}$ (14)
Thus, the correlation between the poisonous gradient baseline $g_{*}^{(t)}$ and the last-iteration aggregated gradient $g_{\Sigma}^{(t-1)}$ is leaked to $S_{2}$, which is $g_{*}^{(t)}=g_{\Sigma}^{(t-1)}+d_{*\Sigma}$ according to Eq. (14).
4.3.3 Gradient Leakage in Clear
Based on the privacy breaches mentioned above, the local gradients can be directly leaked in plaintext, thus breaking the privacy of the whole ShieldFL scheme. Specifically, once $S_{2}$ knows one of the gradients in $\{g_{*}^{(t)},g_{\Sigma}^{(t-1)},\{g_{i}^{(t)}\}_{i\in[1,n]}\}$, it can directly obtain all other users’ local gradients in clear using Eq. (7) or Eq. (13). We present that $S_{2}$ can obtain one of these gradients using the following methods.
  1. a)
    Since $\{g_{*}^{(t)},g_{\Sigma}^{(t-1)},\{g_{i}^{(t)}\}_{i\in[1,n]}\}$ are all normalized gradients, and the difference between any two of these gradients is a known value, it is sufficient to $S_{2}$ to guess a gradient in $\{g_{*}^{(t)},g_{\Sigma}^{(t-1)},\{g_{i}^{(t)}\}_{i\in[1,n]}\}$. For example, $S_{2}$ knows that
    $|\!|g_{j}^{(t)}|\!|={{x}_{j1}^{2}}+\cdots+{x}_{jl}^{2}=1.$ (15)
    According to Eq. (7), $S_{2}$ knows that
    $\tiny\left(\begin{array}[]{c}|\!|g_{1}^{(t)}|\!|\\ \vdots\\ {|\!|g_{i}^{(t)}|\!|}\\ \vdots\\ {|\!|g_{n}^{(t)}|\!|}\\ \end{array}\right)=\left(\begin{array}[]{c}|\!|d_{1j}+g_{j}^{(t)}|\!|\\ \vdots\\ {|\!|d_{ij}+g_{j}^{(t)}|\!|}\\ \vdots\\ {|\!|d_{nj}+g_{j}^{(t)}|\!|}\\ \end{array}\right)=\left(\begin{array}[]{c}\sum_{k=1}^{l}(d_{1j}(k)+x_{jk})^{2}\\ \vdots\\ {\sum_{k=1}^{l}(d_{ij}(k)+x_{jk})^{2}}\\ \vdots\\ {\sum_{k=1}^{l}(d_{nj}(k)+x_{jk})^{2}}\\ \end{array}\right)=\left(\begin{array}[]{c}1\\ \vdots\\ 1\\ \vdots\\ 1\\ \end{array}\right),$ (16)
    $\tiny\begin{split}&\left(\begin{array}[]{c}|\!|g_{\Sigma}^{(t-1)}|\!|\\ \vdots\\ {|\!|g_{\Sigma}^{(t-1)}|\!|}\\ \vdots\\ {|\!|g_{\Sigma}^{(t-1)}|\!|}\\ \end{array}\right)=\left(\begin{array}[]{c}|\!|d_{1j}+g_{j}^{(t)}-d_{1\Sigma}|\!|\\ \vdots\\ {|\!|d_{ij}+g_{j}^{(t)}-d_{i\Sigma}|\!|}\\ \vdots\\ {|\!|d_{nj}+g_{j}^{(t)}-d_{n\Sigma}|\!|}\\ \end{array}\right)\\ =&\left(\begin{array}[]{c}\sum_{k=1}^{l}(d_{1j}(k)-d_{1\Sigma}(k)+x_{jk})^{2}\\ \vdots\\ {\sum_{k=1}^{l}(d_{ij}(k)-d_{i\Sigma}(k)+x_{jk})^{2}}\\ \vdots\\ {\sum_{k=1}^{l}(d_{nj}(k)-d_{n\Sigma}(k)+x_{jk})^{2}}\\ \end{array}\right)=\left(\begin{array}[]{c}1\\ \vdots\\ 1\\ \vdots\\ 1\\ \end{array}\right),\end{split}$ (17)
    where $d_{ij}(k)$ and $d_{i\Sigma}(k)$ are the $k$th element of the vectors $d_{ij}$ and $d_{i\Sigma}$, respectively. Then, if $l\leq 2n+1$, $S_{2}$ can obtain $g_{j}^{(t)}=\{x_{1j},\cdots,x_{lj}\}$ through solving the joint equation of Eq. (15), Eq. (16) and Eq. (17); Otherwise, $S_{2}$ can launch probabilistic attacks to guess $g_{j}^{(t)}$.
  2. b)
    As a practical attack utilized in [35], $S_{2}$ can register as a benign user $\mathcal{U}_{n+1}$ in ShieldFL scheme and submit its local gradients. This behavior does not violate the honest-but-curious assumption of the server in the threat model.

### v2

@@ 4.1 Privacy Leakage in ShieldFL
In ShieldFL, privacy leakages primarily occur during the privacy-preserving defense process. In the procedure of poisonous gradient baseline finding, $S_{1}$ interacts with $S_{2}$ to measure the secure cosine similarity $\cos_{i\Sigma}\leftarrow\mathbf{SecCos}([\![g_{i}^{(t)}]\!],[\![g_{\Sigma}^{(t-1)}]\!])$ for $i\in[1,n]$, so as to find the baseline of poisonous gradient $[\![g_{*}^{(t)}]\!]\in\{[\![g_{i}^{(t)}]\!]\}_{i\in[1,n]}$ with the lowest cosine similarity. Specifically, in this procedure, $S_{1}$ selects $l$ random noises $\{r_{i1},\cdots,r_{il}\}$ and masks $[\![g_{i}^{(t)}]\!]=\{[\![x_{i1}]\!],\cdots,[\![x_{il}]\!]\}$ and $[\![g_{\Sigma}^{(t-1)}]\!]=\{[\![y_{1}]\!],\cdots,[\![y_{l}]\!]\}$ as $[\![\bar{g}_{i}^{(t)}]\!]$ and $[\![\bar{g}_{\Sigma,i}^{(t)}]\!]$ using Eq. (11) and Eq. (22), respectively. Then $S_{1}$ partially decrypts $[\![\bar{g}_{i}^{(t)}]\!]$ and $[\![\bar{g}_{\Sigma,i}^{(t)}]\!]$ as $[\bar{g}_{i}^{(t)}]_{1}$ and $[\bar{g}_{\Sigma,i}^{(t)}]_{1}$ and sends $[\![\bar{g}_{i}^{(t)}]\!],[\![\bar{g}_{\Sigma,i}^{(t)}]\!],[\bar{g}_{i}^{(t)}]_{1},[\bar{g}_{\Sigma,i}^{(t)}]_{1}$ to $S_{2}$.
$\left(\begin{array}[]{c}[\![\bar{g}_{1}^{(t)}]\!]\\ \vdots\\ {[\![\bar{g}_{i}^{(t)}]\!]}\\ \vdots\\ {[\![\bar{g}_{n}^{(t)}]\!]}\\ \end{array}\right)=\left(\begin{array}[]{ccc}[\![x_{11}]\!]\cdot[\![r_{11}]\!]&\cdots&[\![x_{1l}]\!]\cdot[\![r_{1l}]\!]\\ \vdots&\cdot&\vdots\\ {[\![x_{i1}]\!]\cdot[\![r_{i1}]\!]}&\cdots&[\![x_{il}]\!]\cdot[\![r_{il}]\!]\\ \vdots&\cdot&\vdots\\ {[\![x_{n1}]\!]\cdot[\![r_{n1}]\!]}&\cdots&[\![x_{nl}]\!]\cdot[\![r_{nl}]\!]\\ \end{array}\right)$ (11)
@@
$\left(\begin{array}[]{c}[\![\bar{g}_{\Sigma,1}^{(t-1)}]\!]\\ \vdots\\ {[\![\bar{g}_{\Sigma,i}^{(t-1)}]\!]}\\ \vdots\\ {[\![\bar{g}_{\Sigma,n}^{(t-1)}]\!]}\\ \end{array}\right)=\left(\begin{array}[]{ccccc}[\![y_{1}]\!]\cdot[\![r_{11}]\!]&\cdots&[\![y_{l}]\!]\cdot[\![r_{1l}]\!]\\ \vdots&\cdot&\vdots\\ {[\![y_{1}]\!]\cdot[\![r_{i1}]\!]}&\cdots&[\![y_{l}]\!]\cdot[\![r_{il}]\!]\\ \vdots&\cdot&\vdots\\ {[\![y_{1}]\!]\cdot[\![r_{n1}]\!]}&\cdots&[\![y_{l}]\!]\cdot[\![r_{nl}]\!]\\ \end{array}\right)$ (22)
Privacy Leakage: $S_{2}$ decrypts the ciphertexts received from $S_{1}$ to obtain $\bar{g}_{i}^{(t)}$ and $\bar{g}_{\Sigma,i}^{(t-1)}$ $(i\in[1,n])$ as follows:
$\left(\begin{array}[]{c}\bar{g}_{1}^{(t)}\\ \vdots\\ {\bar{g}_{i}^{(t)}}\\ \vdots\\ {\bar{g}_{n}^{(t)}}\\ \end{array}\right)=\left(\begin{array}[]{ccc}x_{11}+r_{11}&\cdots&x_{1l}+r_{1l}\\ \vdots&\cdot&\vdots\\ {x_{i1}+r_{i1}}&\cdots&x_{il}+r_{il}\\ \vdots&\cdot&\vdots\\ {x_{n1}+r_{n1}}&\cdots&x_{nl}+r_{nl}\\ \end{array}\right)$ (33)
@@
$\left(\begin{array}[]{c}\bar{g}_{\Sigma,1}^{(t-1)}\\ \vdots\\ {\bar{g}_{\Sigma,i}^{(t-1)}}\\ \vdots\\ {\bar{g}_{\Sigma,n}^{(t-1)}}\\ \end{array}\right)=\left(\begin{array}[]{ccc}y_{1}+r_{11}&\cdots&y_{l}+r_{1l}\\ \vdots&\cdot&\vdots\\ {y_{1}+r_{i1}}&\cdots&y_{l}+r_{il}\\ \vdots&\cdot&\vdots\\ {y_{1}+r_{n1}}&\cdots&y_{l}+r_{nl}\\ \end{array}\right)$ (44)
@@
Then, $S_{2}$ can learn the difference between two arbitrary local gradients and the difference between the last-iteration aggregated gradient and an arbitrary local gradient:
$\begin{split}d_{ij}&=g_{i}^{(t)}-g_{j}^{(t)}=\{x_{i1}-x_{j1},\cdots,x_{il}-x_{jl}\}\\ &=\left(\bar{g}_{i}^{(t)}-\bar{g}_{j}^{(t)}\right)-\left(\bar{g}_{\Sigma,i}^{(t-1)}-\bar{g}_{\Sigma,j}^{(t-1)}\right),i\neq j.\\ \end{split}$ (45)
@@
The privacy leakage issue in the Byzantine-tolerant aggregation phase is similar to that in the poisonous gradient baseline finding process. For brevity, we present its details in Appendix B. In addition, Appendix B further provides a gradient reconstruction analysis, showing how the leaked information from both phases can be used to recover the complete plaintext gradients.